OpenStudy (anonymous):

Find an explicit rule for the nth term of a geometric sequence where the second and fifth terms are -36 and 2304, respectively.

OpenStudy (anonymous):

do you have a general idea?

OpenStudy (anonymous):

i tried using this http://openstudy.com/study#/updates/513cdcd0e4b029b0182b5689 and i did -36 - ar 2304 = ar^4 2304 = ar * r^3 2304 = -36 * r^3 2304 = r^3????????

OpenStudy (anonymous):

with r being the common ratio but i didn't get an integer

zepdrix (zepdrix):

your second to last step looks correct 2304 = -36 * r^3 what happened to the -36 in the next step? did you mean to divide the 2304 by that amount?

OpenStudy (anonymous):

r^3=2304/-36=?

OpenStudy (anonymous):

-64

OpenStudy (anonymous):

\[r^3=\left( ? \right)^3\] r=?

OpenStudy (anonymous):

cube root it = 4?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

4^3=64

OpenStudy (anonymous):

\[r^3=-64=\left( -4 \right)^3,r=-4\]

OpenStudy (anonymous):

does that mean it would be an = 9 * (-4)^n - 1?

OpenStudy (anonymous):

correct.

OpenStudy (anonymous):

YAY thanks :)))

OpenStudy (anonymous):

yw